# Convergence and Optimality of the EM Algorithm Under Multi-Component Gaussian Mixture Models

### v1

@@ D.1 Proof of Theorem 5: concentration inequalities of the M-steps at the true parameter
$\sup_{v\in\mathbb{S}^{d}}\sum_{a=1}^{L}n_{a}\mathbb{E}\left[\sum_{\ell\in[L]\setminus\{a\}}\gamma_{\ell}(X_{i};\omega^{*})\langle v,X_{i}-\mu_{\ell}^{*}\rangle_{\Sigma^{*}}^{2}\sum_{\ell\in[L]\setminus\{a\}}\gamma_{\ell}(X_{i};\omega^{*})\left\|X_{i}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}^{2}\mid Y_{i}=a\right]$
$\leq 4\sup_{v\in\mathbb{S}^{d}}\sum_{a=1}^{L}n_{a}\sum_{\ell\in[L]\setminus\{a\}}\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})\left((v^{\top}N_{i})^{2}+\Delta_{\max}\right)\left(\|N_{i}\|_{2}^{2}+\Delta_{\max}\right)\mid Y_{i}=a\right]$
$\lesssim n\sum_{a=1}^{L}\pi_{a}^{*}\sum_{\ell\in[L]\setminus\{a\}}{\pi_{\ell}^{*}+\pi_{a}^{*}\over\pi_{a}^{*}}(d+\Delta_{\max})\Delta_{\max}\exp(-c\Delta_{\min})$
$\lesssim nL(d+\Delta_{\max})\Delta_{\max}\exp(-c\Delta_{\min})$ (D.25)
where the penultimate step uses $\mathcal{E}_{\pi}$ as well as Lemma 22. Combining (D.24) and (D.1.3) together with (3.4) yields that
$\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\mathbb{E}[H_{i}^{2}1\{\mathcal{E}_{2,i}\}]\right\|_{\rm op}\lesssim nd+nL(d+\Delta_{\max})\Delta_{\max}\exp(-c\Delta_{\min})\lesssim nd.$
@@
$\textrm{I}_{1}\lesssim\sqrt{dt\over n}+{t(d+\log n+\Delta_{\max})\over n}$ (D.26)
with probability at least $1-de^{-t}-2n^{-1}$. Combined with (D.1.3), the same above holds for I for all $t\geq 1$.
We proceed to bound II. It is worth remarking that a naive application of $\mathcal{E}_{\pi}$ with dual norm inequality leads to a sub-optimal bound of order $\sqrt{L\log n/n}$.
To provide a sharp bound of II, for notational simplicity, we write
$M_{a}:=\mathbb{E}[H_{i}\mid Y_{i}=a],\qquad\forall~a\in[L].$
Note that due to the independence between $Y_{i}$ and $N_{i}$, $M_{a}$’s are simply deterministic matrices. Observing that
II $=\left\|\sum_{a=1}^{L}\left({n_{a}\over n}-\pi_{a}^{*}\right)M_{a}\right\|_{\rm op}={1\over n}\left\|\sum_{i=1}^{n}\left\{\sum_{a=1}^{L}1\{Y_{i}=a\}M_{a}-\mathbb{E}\left[\sum_{a=1}^{L}1\{Y_{i}=a\}M_{a}\right]\right\}\right\|_{\rm op},$
we aim to invoke the matrix-valued Bernstein inequality. To this end, first note that
$\left\|\sum_{a=1}^{L}1\{Y_{i}=a\}M_{a}\right\|_{\rm op}$ $\leq~\max_{a\in[L]}\|M_{a}\|_{\rm op}$
and for any $a\in[L]$,
$\|M_{a}\|_{\rm op}$ $\leq\left\|\mathbb{E}\left[\gamma_{a}(X_{i};\omega^{*})N_{i}N_{i}^{\top}\mid Y_{i}=a\right]\right\|_{\rm op}$
$\quad+\sum_{\ell\in[L]\setminus\{a\}}\left\|\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})\Sigma^{*-1/2}(X_{i}-\mu_{\ell}^{*})(X_{i}-\mu_{\ell}^{*})^{\top}\Sigma^{*-1/2}\mid Y_{i}=a\right]\right\|_{\rm op}$
$\leq 1+2\sum_{\ell\in[L]\setminus\{a\}}\left\|\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})N_{i}N_{i}^{\top}\mid Y_{i}=a\right]\right\|_{\rm op}$
$\quad+2\sum_{\ell\in[L]\setminus\{a\}}\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})\mid Y=a\right]\Delta_{\max}$
$\lesssim 1+\sum_{\ell\in[L]\setminus\{a\}}{\pi_{\ell}^{*}+\pi_{a}^{*}\over\pi_{a}^{*}}\exp(-c\Delta_{\min})$
$\lesssim 1$ (D.27)
by invoking Lemma 22 in the penultimate step and using (3.4) in the last step. Furthermore,
$\left\|\sum_{i=1}^{n}\mathbb{E}\left[\left(\sum_{a=1}^{L}1\{Y_{i}=a\}M_{a}\right)^{2}\right]\right\|_{\rm op}$ $=n\left\|\sum_{a=1}^{L}\pi_{a}^{*}M_{a}^{2}\right\|_{\rm op}\leq n\|M_{a}\|_{\rm op}\|\sum_{a=1}^{L}\pi_{a}^{*}M_{a}\|_{\rm op}\lesssim n$
by using (D.21) and (D.27). Applying Lemma 35 with $t=C\sqrt{n\log n}$ gives that
$\mathbb{P}\left\{\textrm{II}\lesssim\sqrt{\log n\over n}\right\}\geq 1-Ln^{-2}\geq 1-n^{-1}.$ (D.28)
Since by (D.12), with probability at least $1-3n^{-1}$,
III $\lesssim\sum_{\ell=1}^{L}\pi_{\ell}^{*}\left\|\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}^{2}\lesssim{dL\log n\over n}+{L\log^{2}n\over n^{2}\pi_{\min}^{*}}\Delta_{\max}$ (D.29)
By collecting Eqs. D.28, D.29 and D.26 with $t=2\log n$ and using $d\leq n$ and $\log n\leq n\pi_{\min}^{*}\leq n/L$, we conclude that with probability at least 14/n1-4/n,
$d(\bm{\widehat{\Sigma}}(\omega^{*}),\bm{\Sigma}(\omega^{*}))$ $\lesssim\sqrt{d\log n\over n}+{dL\log n\over n}+\left(1+{L\log n\over n\pi_{\min}^{*}}\right){\log n\over n}\Delta_{\max}$
Using $\Delta_{\min}\leq C\log n$ thus gives (4.4).
Case 2: $\Delta_{\min}\geq C\log n$.
Based on (D.1.3), we bound $d(\bm{\widehat{\Sigma}}(\omega^{*}),\bm{\Sigma}(\omega^{*}))\leq\textrm{I}^{\prime}+\textrm{II}^{\prime}+\textrm{III}$ with III given by (D.22) and
$\textrm{I}^{\prime}$ $={1\over n}\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\Sigma^{*-1/2}\gamma_{a}(X_{i};\omega^{*})(X_{i}-\mu_{a}^{*})(X_{i}-\mu_{a}^{*})^{\top}\Sigma^{*-1/2}-\mathbf{I}_{d}\right\|_{\rm op}$
$\textrm{II}^{\prime}$ $={1\over n}\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\sum_{\ell\in[L]\setminus\{a\}}\Sigma^{*-1/2}\gamma_{\ell}(X_{i};\omega^{*})(X_{i}-\mu_{\ell}^{*})(X_{i}-\mu_{\ell}^{*})^{\top}\Sigma^{*-1/2}\right\|_{\rm op}.$
We work on the event $\mathcal{E}_{\Delta}$ given by (D.14). First,
$\textrm{I}^{\prime}$ $\leq{1\over n}\left\|\sum_{i=1}^{n}N_{i}N_{i}^{\top}-\mathbf{I}_{d}\right\|_{\rm op}+{1\over n}\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}(1-\gamma_{a}(X_{i};\omega^{*}))N_{i}N_{i}^{\top}\right\|_{\rm op}$
[…]
Furthermore, by (D.17),
$\textrm{II}^{\prime}$ $\leq{3\over n}\left\|\sum_{a=1}^{L}\sum_{i\in\widehat{G}_{a}}\sum_{\ell\in[L]\setminus\{a\}}{\pi_{\ell}^{*}\over\pi_{a}^{*}}\left(N_{i}N_{i}^{\top}+\Sigma^{*-1/2}(\mu_{\ell}^{*}-\mu_{a}^{*})(\mu_{\ell}^{*}-\mu_{a}^{*})^{\top}\Sigma^{*-1/2}\right)\right\|_{\rm op}\exp(-\Delta_{\min}/4)$
$\leq 3\left({1\over n}\left\|\sum_{i=1}^{n}N_{i}N_{i}^{\top}\right\|_{\rm op}+\Delta_{\max}\right){1\over\pi_{\min}^{*}}\exp(-\Delta_{\min}/4).$
Also note that on $\mathcal{E}_{\pi}$,
$\textrm{III}\lesssim\sum_{\ell=1}^{L}\pi_{\ell}^{*}\left\|\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}^{2}.$
By invoking Lemma 33, Lemma 36 and (D.19) together with $d\log n\leq n$, we obtain that
$d(\bm{\widehat{\Sigma}}(\omega^{*}),\bm{\Sigma}(\omega^{*}))$ $\lesssim{\Delta_{\max}\over\pi_{\min}^{*}}\exp(-\Delta_{\min}/4)+\sqrt{d+\log n\over n}+{(d+\log n)L\over n}$
$\lesssim\sqrt{d+\log n\over n}+{(d+\log n)L\over n}$
with probability $1-3n^{-1}$. This completes the proof of Case 2, hence concluding the bound in (4.4). ∎
@@ D.2 Contraction of the sample level M-steps
$\nabla^{\pi}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})$ $:=\sum_{k=1}^{L-1}\int_{0}^{1}\frac{1}{\widehat{\bm{\pi}}_{\ell}(\omega)}\mathbb{E}_{n}\left[\left(X-\widehat{\bm{\mu}}_{\ell}(\omega)\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\pi_{k}}\Big{|}_{\omega=(\pi_{u},\bar{M},\bar{J})},~\left(\bar{\pi}_{k}-\pi_{k}^{*}\right)\right\rangle\right]{\rm d}u$
$\nabla^{M}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})$ $:=\sum_{k=1}^{L}\int_{0}^{1}\frac{1}{\widehat{\bm{\pi}}_{\ell}(\omega)}\mathbb{E}_{n}\left[\left(X-\widehat{\bm{\mu}}_{\ell}(\omega)\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\mu_{k}}\Big{|}_{\omega=(\pi^{*},M_{u},\bar{J})},~\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle\right]{\rm d}u$
$\nabla^{J}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})$ $:=\int_{0}^{1}\frac{1}{\widehat{\bm{\pi}}_{\ell}(\omega)}\mathbb{E}_{n}\left[\left(X-\widehat{\bm{\mu}}_{\ell}(\omega)\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial J}\Big{|}_{\omega=(\pi^{*},M^{*},J_{u})},~\bar{J}-J^{*}\right\rangle\right]{\rm d}u.$
Analogous to (C.43), we define
$\widehat{D}_{\mu_{\ell}}:=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\|\widehat{\bm{\mu}}_{\ell}(\omega)-\bm{\mu}_{\ell}(\omega^{*})\|_{\Sigma^{*}}.$ (D.37)
@@
By similar arguments of proving Eqs. C.49, C.51 and C.53, we find that
$\left\|\nabla^{\pi}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left[\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(1)}}+\widehat{\mathcal{R}}_{I}^{(0)}\left(\sqrt{\Delta_{\max}}+\widehat{D}_{\mu_{\ell}}\right)\right]d(\bar{\pi},\pi^{*})$
$\left\|\nabla^{M}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left[\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(1)}}+\widehat{\mathcal{R}}_{I}^{(0)}\sqrt{\Delta_{\max}}\left(\sqrt{\Delta_{\max}}+\widehat{D}_{\mu_{\ell}}\right)\right]d(\bar{M},M^{*})$
$\left\|\nabla^{J}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left[\sqrt{\widehat{\mathcal{R}}_{I}^{(1)}\widehat{\mathcal{R}}_{I}^{(2)}}+\left(\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}+\widehat{\mathcal{R}}_{I}^{(0)}\sqrt{\Delta_{\max}}\right)\left(\sqrt{\Delta_{\max}}+\widehat{D}_{\mu_{\ell}}\right)\right]d(\bar{J},J^{*}).$
@@
$\left\|\widehat{\bm{\mu}}_{\ell}(\bar{\omega})-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}$~ $\leq~\sqrt{\widehat{\mathcal{R}}_{I}^{(1)}\widehat{\mathcal{R}}_{I}^{(2)}}d(\bar{J},J^{*})+\left(\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}+\widehat{\mathcal{R}}^{(0)}_{I}\sqrt{\Delta_{\max}}\right)d(\bar{\omega},\omega^{*})$
$\quad+\left[\widehat{\mathcal{R}}^{(0)}_{I}d(\bar{\omega},\omega^{*})+\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}d(\bar{J},J^{*})\right]\widehat{D}_{\mu_{\ell}}.$
On the event (D.32), it implies
$\widehat{D}_{\mu_{\ell}}\lesssim\sqrt{\widehat{\mathcal{R}}_{I}^{(1)}\widehat{\mathcal{R}}_{I}^{(2)}}\sqrt{\Delta_{\min}}+\left(\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}+\widehat{\mathcal{R}}^{(0)}_{I}\sqrt{\Delta_{\max}}\right)\sqrt{\Delta_{\max}\Delta_{\min}}\lesssim\sqrt{\Delta_{\max}}$
so that
$\left\|\widehat{\bm{\mu}}_{\ell}(\bar{\omega})-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}$~ $\lesssim~\sqrt{\widehat{\mathcal{R}}_{I}^{(1)}\widehat{\mathcal{R}}_{I}^{(2)}}d(\bar{J},J^{*})+\left(\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}+\widehat{\mathcal{R}}^{(0)}_{I}\sqrt{\Delta_{\max}}\right)d(\bar{\omega},\omega^{*}).$
@@ D.2.2 An $\epsilon$-net construction of $\mathcal{B}_{d}(\omega^{*})$
Finally, regarding $\pi$, there exists an $\epsilon_{3}$-net $\mathcal{N}_{\epsilon_{3}}(\mathcal{S}^{L})$ of the simplex in $\mathbb{R}^{L}$ with respect to the sup-norm [19][Lemma A.4]. As a result, for any $\pi$ such that $d(\pi,\pi^{*})\leq c_{0}$, there exists $\pi^{\prime}\in\mathcal{N}_{\epsilon_{3}}(\mathcal{S}^{L})$ such that

### v2

@@ D.1 Proof of Theorem 5: concentration inequalities of the M-steps at the true parameter
$\sup_{v\in\mathbb{S}^{d}}\sum_{a=1}^{L}\pi_{a}^{*}\mathbb{E}\left[\sum_{\ell\in[L]\setminus\{a\}}\gamma_{\ell}(X_{i};\omega^{*})\langle v,X_{i}-\mu_{\ell}^{*}\rangle_{\Sigma^{*}}^{2}\sum_{\ell\in[L]\setminus\{a\}}\gamma_{\ell}(X_{i};\omega^{*})\left\|X_{i}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}^{2}\mid Y_{i}=a\right]$
$\leq 4\sup_{v\in\mathbb{S}^{d}}\sum_{a=1}^{L}\pi_{a}^{*}\sum_{\ell\in[L]\setminus\{a\}}\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})\left((v^{\top}N_{i})^{2}+\Delta_{a\ell}\right)\left(\|N_{i}\|_{2}^{2}+\Delta_{a\ell}\right)\mid Y_{i}=a\right]$
$\lesssim\sum_{a=1}^{L}\sum_{\ell\in[L]\setminus\{a\}}(\pi_{\ell}^{*}+\pi_{a}^{*})(d+\Delta_{a\ell})\Delta_{a\ell}\exp(-c\Delta_{a\ell})$
$\lesssim dL\exp(-c\Delta_{\min})$ (D.13)
where the penultimate step uses Lemma 21. Combining (D.12) and (D.1.3) together with (3.4) yields that
$\left\|\sum_{i=1}^{n}\mathbb{E}[H_{i}^{2}1\{\mathcal{E}_{2,i}\}]\right\|_{\rm op}\lesssim nd+ndL\exp(-c\Delta_{\min})\lesssim nd.$
@@
$\textrm{I}_{1}\lesssim\sqrt{dt\over n}+{tL(d+\log n)\over n}$ (D.14)
with probability at least $1-de^{-t}-2n^{-1}$. Regarding $\textrm{I}_{2}$, by using the arguments in (D.12) and (D.1.3), one can deduce that
$\textrm{I}_{2}\leq\max_{i\in[n]}\left\|\mathbb{E}\left[H_{i}1\{\mathcal{E}_{2,i}^{c}\}\right]\right\|_{{\rm op}}\leq\max_{i\in[n]}\sqrt{\|\mathbb{E}[H_{i}\|H_{i}\|_{\rm op}]\|_{\rm op}}\sqrt{\mathbb{P}(\mathcal{E}_{2,i}^{c})}\lesssim{\sqrt{d}\over n}.$
Thus the bound in (D.14) holds for I for all $t\geq 1$. In conjunction with (D.11), by choosing $t=2\log n$ and using $d\leq n$ and $\log n\leq n\pi_{\min}^{*}\leq n/L$, we conclude that with probability at least 14/n1-4/n,
$d(\bm{\widehat{\Sigma}}(\omega^{*}),\bm{\Sigma}(\omega^{*}))$ $\lesssim\sqrt{d\log n\over n}+{(d+\log n)L\log n\over n}$
completing the proof of (4.4). ∎
@@ D.2 Contraction of the sample level M-steps
$\nabla^{\pi}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})$ $:=\sum_{k=1}^{L-1}\int_{0}^{1}\frac{1}{\widehat{\bm{\pi}}_{\ell}(\omega)}\mathbb{E}_{n}\left[\left(X-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\pi_{k}}\Big|_{\omega=(\pi_{u},\bar{M},\bar{J})},~\left(\bar{\pi}_{k}-\pi_{k}^{*}\right)\right\rangle\right]{\rm d}u$
$\nabla^{M}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})$ $:=\sum_{k=1}^{L}\int_{0}^{1}\frac{1}{\widehat{\bm{\pi}}_{\ell}(\omega)}\mathbb{E}_{n}\left[\left(X-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\mu_{k}}\Big|_{\omega=(\pi^{*},M_{u},\bar{J})},~\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle\right]{\rm d}u$
$\nabla^{J}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})$ $:=\int_{0}^{1}\frac{1}{\widehat{\bm{\pi}}_{\ell}(\omega)}\mathbb{E}_{n}\left[\left(X-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial J}\Big|_{\omega=(\pi^{*},M^{*},J_{u})},~\bar{J}-J^{*}\right\rangle\right]{\rm d}u.$
@@
Meanwhile the event (D.20) implies
$\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})-\mu_{b}^{*}\|_{\Sigma^{*}}\leq\sqrt{\Delta_{\max}}+d(\widehat{\mathbf{M}}(\omega^{*}),\mathbf{M}(\omega^{*}))\leq 2\sqrt{\Delta_{\max}}.$ (D.22)
By similar arguments of proving Eqs. C.47, C.49 and C.51, we find that
$\left\|\nabla^{\pi}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left[\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(1)}}+\widehat{\mathcal{R}}_{I}^{(0)}\sqrt{\Delta_{\max}}\right]d(\bar{\pi},\pi^{*})$
$\left\|\nabla^{M}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left[\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(1)}}+\widehat{\mathcal{R}}_{I}^{(0)}\Delta_{\max}\right]d(\bar{M},M^{*})$
$\left\|\nabla^{J}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left[\sqrt{\widehat{\mathcal{R}}_{I}^{(1)}\widehat{\mathcal{R}}_{I}^{(2)}}+\left(\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}+\widehat{\mathcal{R}}_{I}^{(0)}\sqrt{\Delta_{\max}}\right)\sqrt{\Delta_{\max}}\right]d(\bar{J},J^{*}).$
@@
$\left\|\widehat{\bm{\mu}}_{\ell}(\bar{\omega})-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}$~ $\leq~\sqrt{\widehat{\mathcal{R}}_{I}^{(1)}\widehat{\mathcal{R}}_{I}^{(2)}}d(\bar{J},J^{*})+\left(\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}+\widehat{\mathcal{R}}^{(0)}_{I}\sqrt{\Delta_{\max}}\right)d(\bar{\omega},\omega^{*}),$
as desired.
@@ D.2.2 An $\epsilon$-net construction of $\mathcal{B}_{d}(\omega^{*})$
Finally, regarding $\pi$, there exists an $\epsilon_{3}$-net $\mathcal{N}_{\epsilon_{3}}(\mathcal{S}^{L})$ of the simplex in $\mathbb{R}^{L}$ with respect to the sup-norm [ghosal2001entropies][Lemma A.4]. As a result, for any $\pi$ such that $d(\pi,\pi^{*})\leq c_{0}$, there exists $\pi^{\prime}\in\mathcal{N}_{\epsilon_{3}}(\mathcal{S}^{L})$ such that